# Human-Object Interaction from Human-Level Instructions

### v2

@@ 1 Introduction
The generated motion, however, is not guaranteed to be physically accurate. To ensure physical plausibility, we further use reinforcement learning (RL) to track the generated motion in a physics simulator. Our method supports tracking diverse interactions with multiple distinct objects within a single sequence.
@@ 2 Related Work
There has been increasing attention on synthesizing human-object interactions with dynamic objects [6, 35, 55, 62]. OMOMO [25] collected a dataset for dynamic human-object interactions and proposed a framework to synthesize human motions from object motions. InterDiff [61] predicts future interaction motions based on past object and human motion. CHOIS [24] synthesizes interactions in 3D scenes from text and sparse waypoints. Despite these advancements, all of these works focus on full-body human motion generation without detailed finger motions. In this work, we aim to generate synchronized object motions, human motions, and finger motions.
@@ 5.1 Background: Conditional Diffusion Model
where $p_{\theta}$ represents a neural network, $\bm{c}$ represents the conditions, $\bm{\mu}_{\theta}(\bm{x}_{n},n,\bm{c})$ represents the learned mean, $\sigma_{n}^{2}$ denotes a fixed variance. The training loss is defined as $\mathcal{L}=\mathbb{E}_{\bm{x}_{0},n}||\hat{\bm{x}}_{\theta}(\bm{x}_{n},n,\bm{c})-\bm{x}_{0}||_{1}$.
@@ 5.2.1 CoarseNet
Given the initial human and object poses, waypoints, and final object pose, along with a text instruction, CoarseNet generates synchronized human motion and object motion. This is achieved using a pre-trained diffusion model [24].
@@ 5.2.2 Grasp Pose Generation
The optimized grasp pose is denoted by $\hat{\bm{g}}=(\hat{\bm{w}},{\hat{\bm{\theta}}})$, where $\hat{\bm{w}}$ is the optimized wrist pose, and $\hat{\bm{\theta}}$ is the finger pose. This grasp pose $\hat{\bm{g}}$ is maintained throughout the contact phase. However, as CoarseNet’s generated motion may not fully align with $\hat{\bm{w}}$, we introduce RefineNet to address this misalignment.
@@ 7 Experiment
We evaluate each system module by comparing our proposed approach with baseline methods and conducting ablation studies. We highly recommend that readers refer to the supplementary video for additional qualitative results.
@@ 7.1 Evaluation of High-Level Planner
The core component of the high-level planner is the scene map calculation. To assess its effectiveness, we compare our method to a baseline that uses LLMs to directly predict object positions and orientations. We design 15 human-level instructions covering common everyday tasks for assessment, using OpenAI’s GPT-4o [33] for both methods.
Figure 4: Comparison between (a) the baseline and (b) our LLM planner. The baseline generates an incorrect monitor position and orientation, and places the chair in a position that intersects with the table. The human perception study results (bottom) show that the majority of participants preferred our results.
@@
Figure 5: Qualitative comparisons of each method. All baselines and ablations cannot produce precise hand and finger motions, while our method generates natural hand-object interactions.
We conduct a human perception study to evaluate the methods, with results presented in Fig. 4. Each pair was reviewed by 20 Amazon Mechanical Turk (AMT) workers. As shown in Fig. 4, the baseline method fails to correctly position the objects and accurately determine their facing directions.
@@ 7.2 Evaluation of Low-Level Motion Generator
Following prior work [24], we evaluate our approach from various aspects. For condition matching, we report the waypoint following errors ($T_{xy}$) in centimeters. For human motion quality, we report the foot sliding score (FS) and foot height ($H_{\text{feet}}$), both measured in centimeters. For interaction quality, we use contact precision (Cprec), recall (Crec), and F1 score (CF1). To better assess finger-object interaction, we also report contact coverage (CC) – the percentage of hand points within -2mm to +2mm of the object surface – and intersection volume (IV), the overlap volume of hand and object meshes in $\text{cm}^{3}$, following [11]. For ground truth difference, we report the mean per-joint position error (MPJPE), object position error ($T_{\text{obj}}$), both in centimeters (cm), along with the object orientation error ($O_{\text{obj}}$). We also conduct a human perception study to measure the motion quality.
@@
For the task of generating full-body human motion, finger motion, and object motion from text, prior work [9, 26] trained models on GRAB [43] to predict motions that manipulate given objects. GRAB [43] contains small object manipulation data without locomotion. Models trained on GRAB cannot generalize to scenarios involving large-sized objects that require full-body coordination and locomotion. The approach from [9, 26] cannot be directly applied to train on the FullBodyManipulation dataset [25] since this dataset does not contain finger motions, making it infeasible to learn a model via direct supervision. Thus, we establish our baseline by combining two complementary methods: CHOIS [24], which generates object motion and full-body motion from text, and GRIP [45], which refines arm movements and generates finger motions based on CHOIS’s output.
In addition, to evaluate the impact of each component, we compare the full system to two ablations: (1) CoarseNet alone (CNet) and (2) CoarseNet + RefineNet (C+RNet). This evaluation focuses on the motion synthesis quality before physics tracking. We present the evaluation for the physics tracker in the next subsection.
@@ Results.
The qualitative results are shown in Fig. 5. Both ablations exhibit a lack of finger details, and GRIP [45] fails to generate accurate hand motion, leading to significant artifacts. In contrast, our method generates precise finger-object interactions. The quantitative results in Tab. 1 demonstrate our method’s superior performance in terms of interaction quality. We also observe the impact of RefineNet: compared to CNet, C+RNet achieves better interaction quality. As shown in Fig. 5, C+RNet maintains reasonable relative poses between the hand and object throughout the interaction, while CNet exhibits more artifacts.
@@ 7.3 Evaluation of Physics Tracker
The physics tracker can accurately track kinematic motion with an average tracking error of approximately 2 cm per frame per joint. Tracking error is defined as the position error between the reference motion and the tracked motion. Through physics simulation, artifacts such as foot floating and human-object penetration are eliminated, and foot sliding is also significantly reduced, with an FS metric of 0.002. Qualitative results are shown in Fig. 7.
We also compare our approach with PhysHOI [52]. While [52] demonstrates impressive results in basketball playing motion, it does not perform well in our task. The policy fails to grasp objects, as shown in Fig. 8.
Figure 7: Qualitative comparison of kinematic motion and physics-based motion. Our physics tracker corrects artifacts such as foot floating and penetration.
Figure 8: PhysHOI [52] fails to track kinematic motion, while our policy successfully drives the character to interact with diverse objects.
@@ 8 Conclusions
In this work, we introduced a system that simulates realistic human-object manipulation in a contextual environment based on human-level language instructions. The framework consists of a high-level LLM planner that infers a target scene layout and execution plan from the given instructions, a low-level motion generator that generates synchronized, full-body, fingers, and object motion, and a physics tracker that imitate the generated kinematic motions and produces physically plausible interaction motions. Our system effectively handles a variety of tasks, showcasing its potential for various real-world applications.
Limitations.
Our system currently uses text or image inputs to represent the contextual environment, and could potentially be enhanced with more structured representations like voxel grids as used in [20]. While our approach successfully leverages LLMs to interpret human-level instructions, there is room to improve plan generation for more complex environments. For grasp pose generation, we build upon [50], though alternative approaches could be explored to further improve grasp success rates.
@@
A promising direction is creating an agent with egocentric vision perception as input to LLM/VLMs for high-level motion planning. This could enable the agent to exhibit more natural behaviors, such as purposeful head orientation for observation and better body positioning to avoid collisions.

### v3

@@ 1 Introduction
The generated motion, however, is not guaranteed to be physically accurate. To ensure physical plausibility, we further use reinforcement learning (RL) to track the generated motion in a physics simulator. Our method adopts an importance sampling technique to improve the training efficiency of long sequence tracking and supports tracking diverse interactions with multiple distinct objects within a single sequence.
@@ 2 Related Work
There has been increasing attention on synthesizing human-object interactions with dynamic objects [6, 38, 58, 64]. OMOMO [26] collected a dataset for dynamic human-object interactions and proposed a framework to synthesize human motions from object motions. CHOIS [25] synthesizes interactions in 3D scenes from text and sparse waypoints. Despite these advancements, all of these works focus on full-body human motion generation without detailed finger motions. In this work, we aim to generate synchronized object motions, human motions, and finger motions.
@@ 5.1 Background: Conditional Diffusion Model
[…]
@@ 5.2.1 CoarseNet
Given the initial human and object poses, waypoints, and final object pose, along with a text instruction, CoarseNet generates synchronized human motion and object motion. This is achieved using a pre-trained diffusion model CHOIS [25].
@@ 5.2.2 Grasp Pose Generation
[…]
@@ 7 Experiment
We conduct extensive evaluation to assess the overall performance of our system, as well as each individual component. We use quantitative metrics and user perception studies to evaluate our results compared to the baselines. We also conduct ablation studies to analyze the effects of each system module. We present the key results in this section and highly recommend readers refer to the supplementary video and documents for more detailed results and descriptions of the studies.
@@ 7.1 Evaluation of High-Level Planner
The core component of the high-level planner is the scene map calculation. To assess its effectiveness, we quantitatively measure the geometric accuracy of the resulting plans and conduct a human perception study. For evaluation, we compare our method to a baseline that uses LLMs to directly predict object positions and orientations. We design 25 human-level instructions covering common everyday tasks for assessment, using OpenAI’s GPT-4o [35] for both our method and the baseline.
Figure 4: Comparison between (a) the baseline and (b) our high-level planner. The baseline generates an incorrect monitor position and orientation, and places the chair in a position that intersects with the table. The human perception study results (bottom) show that the majority of participants preferred our results.
Method Condition Human Motion Interaction GT Difference $T_{xy}\downarrow$ $H_{\text{feet}}\downarrow$ FS $\downarrow$ $\text{C}_{\text{prec}}\uparrow$ $\text{C}_{\text{rec}}\uparrow$ $\text{C}_{\text{F1}}\uparrow$ $C$ $\uparrow$ IV $\downarrow$ $\text{MPJPE}\downarrow$ $T_{\text{obj}}\downarrow$ $O_{\text{obj}}\downarrow$ CNet+GRIP [48] 3.58 2.50 0.45 0.90 0.64 0.70 3.99% 49.00 14.33 11.55 0.93 CNet 3.26 2.68 0.43 0.91 0.81 0.84 3.54% 48.56 14.32 10.56 1.06 C+RNet 2.81 2.89 0.33 0.91 0.95 0.92 3.84% 24.78 15.96 12.57 0.73 C+R+FNet (ours) 2.81 2.89 0.33 0.91 0.95 0.92 5.53% 19.06 15.96 12.57 0.73
Figure 5: Qualitative comparisons of each method. All baselines and ablations fail to produce precise hand and finger motions, while our method generates natural hand-object interactions.
Table 2: Comparison of geometric accuracy between our method and the baseline.
$_{\text{p}}\downarrow PE$$_{\text{o}}\downarrow$ Baseline 21.9% 12.5% Ours 3.1% 1.6%
@@
We also conducted a human perception study, with results presented in the lower part of Fig. 4. Each pair was reviewed by 20 Amazon Mechanical Turk (AMT) workers. They were asked to choose the scene that was more reasonable and consistent with the given instruction. The majority of participants preferred our method.
@@ 7.2 Evaluation of Low-Level Motion Generator
Following prior work [25], we evaluate our approach from various aspects. All distance errors are measured in centimeters. For condition matching, we report the waypoint following errors ($T_{xy}$). For human motion quality, we report the foot sliding score (FS) and foot height ($H_{\text{feet}}$). For interaction quality, we use contact precision (Cprec), recall (Crec), and F1 score (CF1). To better assess finger-object interaction, we also report contact coverage (CC) – the percentage of hand points within -2mm to +2mm of the object surface – and intersection volume (IV), the overlap volume of hand and object meshes in $\text{cm}^{3}$, following [11]. For ground truth difference, we report the mean per-joint position error (MPJPE), object position error ($T_{\text{obj}}$), and the object orientation error ($O_{\text{obj}}$). We also conduct a human perception study to measure the motion quality.
@@
For generating full-body human motion, finger motion, and object motion from text, prior work [9, 27] trained models on GRAB [46] to predict motions for small-object manipulation. However, GRAB contains only small-object manipulation data without locomotion, so models trained on it cannot generalize to scenarios involving large objects that require full-body coordination and locomotion. Moreover, the approach from [9, 27] is not applicable to the FullBodyManipulation dataset [26] used in our work, as this dataset lacks finger motions, making it infeasible to learn a model via direct supervision. Due to these differences, a directly comparable baseline does not exist. Thus, we establish our baseline by combining two complementary methods: CHOIS [25] (referred to as CoarseNet in our framework), which generates object motion and full-body motion from text, and GRIP [48], which refines arm movements and generates finger motions based on CHOIS’s output.
In addition, to evaluate the impact of each component, we compare our full system against two ablations: (1) CoarseNet alone (CNet) and (2) CoarseNet + RefineNet (C+RNet), which represents the full system without finger motion. The full system is denoted as C+R+FNet. This evaluation focuses on the motion synthesis quality before physics tracking. We present the evaluation for the physics tracker in Sec. 7.3.
@@ Results.
The qualitative comparison in Fig. 5 clearly demonstrates the superiority of our method. Both ablated variants show a lack of finger detail, and GRIP [48] fails to generate accurate hand motion, resulting in significant artifacts. In contrast, our method produces precise finger-object interactions. We also observe the impact of RefineNet: compared to CNet, C+RNet maintains more consistent and reasonable relative poses between the hand and object during the interaction, showing RefineNet’s role in improving pose consistency and reducing artifacts.
The quantitative results are presented in Tab. 1. Since our method focuses on finger-object interaction, the five metrics under the Interaction column are particularly important, and we achieve the best performance across all of them.
@@ 7.3 Evaluation of Physics Tracker
The physics tracker can accurately track kinematic motion while ensuring the physical plausibility of the motion. Through physics simulation, artifacts such as foot floating, foot sliding, and human-object penetration are eliminated, as shown in Fig. 7.
Table 3: Tracking error between the kinematic motion and physics-based motion.
$E_{h}$ $E_{o}$ Ours 5.45 4.67
Additionally, we evaluate the tracking error of human joints ($E_{h}$) and objects ($E_{o}$), as shown in Tab. 3. The tracking error is defined as the positional error (cm) between the kinematic motion and the tracked motion. The results were evaluated across 20 interaction sequences, each with an average duration of 30 seconds and 2 objects of varying shapes and types. All sequences were successfully tracked, with relatively small errors, demonstrating high tracking accuracy.
Figure 7: Qualitative comparison of kinematic motion and physics-based motion. The physics tracker corrects artifacts such as foot floating and penetration.
@@ 8 Conclusions
In this work, we introduced a system that simulates realistic human-object manipulation in a contextual environment based on human-level language instructions. The framework consists of a high-level LLM planner that infers a target scene layout and execution plan from the given instructions, a low-level motion generator that generates synchronized, full-body, fingers, and object motion, and a physics tracker that imitates the generated kinematic motions and produces physically plausible interaction motions. Our system effectively handles a variety of tasks, showcasing its potential for various real-world applications.
@@
Our system currently uses text or image inputs to represent the contextual environment, and could potentially be enhanced with more structured representations like voxel grids as used in [20]. Building on this idea, a promising direction is to create an agent with egocentric vision perception as input to LLM/VLMs for high-level motion planning. This could enable the agent to exhibit more natural behaviors, such as purposeful head orientation for observation and better body positioning to avoid collisions.
@@ Acknowledgement.
Supplementary Material
A Details of High-Level Planner
A.1 Pseudo-code for Object Positions Calculation
In Algorithm 1, we provide the pseudo-code for the algorithm that calculates object positions in the scene graph, as described in Sec. 4.1.
Input : Scene graph $G(V,E)$ with vertices $V$ and edges $E$, initial positions $L$, static nodes ($V_{S}$), the nodes to be moved ($V_{M}$).
Output : Updated positions $L$ for all nodes.
Procedure ComputePositions($G$, $L$):
  Let the set of processed nodes $V_{P}\leftarrow V_{S}$
 while $V_{M}$ is not empty do
     Let $V^{\prime}\subseteq V_{M}$ be the nodes whose predecessors’ positions are known
     foreach $v\in V^{\prime}$ do
        $L(v)\leftarrow$ Update($v$, $G$, $L$)
        $V_{P}\leftarrow V_{P}\cup\{v\}$
        $V_{M}\leftarrow V_{M}\setminus\{v\}$
     end foreach
  end while
 return $L$
return
Procedure Update($v$, $G$, $L$):
  Initialize an empty list of positions $P$
  foreach predecessor $u$ of $v$ in $G$ do
     $O\leftarrow$ ComputeOffset($(u,v)$, $L(u)$)
     Append $L(u)+O$ to $P$
  end foreach
 return $\text{Average}(P)$
return
Procedure ComputeOffset($(u,v)$, $L(u)$):
  if edge $(u,v)$ indicates “on” then
     Set $O_{\text{height}}$ to align with the top surface of $u$
     Sample $O_{\text{horizontal}}$ within the horizontal bounds of $u$
     return $O_{\text{horizontal}},O_{\text{height}}]$
  end if
 else if edge $(u,v)$ indicates “adjacent” then
     Use the direction and distance provided by the LLMs to compute $O$
     return $O$
  end if
return
ALGORITHM 1 Object Positions Calculation
A.2 Prompts for High-Level Planner
We provide the prompts used for generating the scene map and execution plan in Fig. S1. An example of the input and the corresponding output of the LLM planner is shown in Fig. S2. We also present the prompt used by the baseline (as described in Sec. 7.1) in Fig. S3.
Figure S1: Prompts used by the LLM planner to generate the scene map and execution plan.
Figure S2: Example of the input and output for the LLM planner.
Figure S3: Prompts used by the baseline, as described in Sec. 7.1.
B Details of Low-Level Motion Generator
B.1 Diffusion Model Architecture
Our system includes four diffusion models: CoarseNet, RefineNet, FingerNet, and the navigation module, all of which adopt a similar transformer-based architecture. Here, we illustrate the architecture of RefineNet in Fig. S4. The model is provided with a motion sequence $\bm{x}_{n}$, the condition $\bm{c}_{r}$, and the noise step $n$, and predicts the clean motion $\hat{\bm{x}}_{0}$. For the other networks, we simply substitute the condition with the corresponding one.
Figure S4: Architecture of RefineNet. The model is conditioned on $\bm{c}_{r}=\left\{\bm{W},\bm{S}_{r},\bm{G},\bm{T}\right\}$, where $\bm{W}$ represents the wrist-object relative pose, $\bm{S}_{r}$ is the masked motion representation containing the static pose and waypoints, $\bm{G}$ denotes the object geometry, and $\bm{T}$ is the textual input.
B.2 Implementation Details of Grasp Pose Generation
As introduced in Sec. 5.2.2, we use a state-of-the-art grasp pose generation method  [53] to generate the grasp pose. Although originally designed for robotic hand grasping, this method relies solely on kinematic chain information, making it fully compatible with the SMPL-X [37] hand model. Starting with an object mesh and an initial hand pose, the grasp generation process is formulated as an optimization problem that minimizes a differentiable force closure energy term along with several additional terms that encourage the fingers to conform closely to the object surface while maintaining a natural hand shape. For motions involving both hands, such as carrying a box, we omit the force closure term.
B.3 Post-processing in RefineNet
As introduced in Sec. 5.2.3, we apply post-processing to RefineNet’s output to better align it with the input condition. First, we replace the object poses in the pre-contact and post-contact phases with the corresponding static poses. To address the abrupt transitions at the phase boundaries, we identify the difference between the predicted object pose and the static object pose at the boundary frame denoted as $\Delta\bm{d}\in\mathbb{R}^{6}$ (suppose this is the boundary between pre-contact phase and contact phase), which encapsulates both the positional and rotational differences in axis-angle format. To smooth these transitions, we employ an interpolation strategy to the object motion in the contact phase. The interpolation function is defined by $\hat{\bm{O}}_{t}=\bm{O}_{t}+\alpha_{t}\Delta\bm{d}$, where $\alpha_{t}$ is a time-dependent scaling factor that decreases from 1 to 0 across the transition period, $\bm{O}_{t}$ denotes the predicted object pose.
[…]
B.4 Implementation Details of FingerNet
FingerNet generates finger motions for the pre-contact and post-contact phases to create natural transitions between rest and grasp poses. It is trained on the GRAB dataset [46]. Since FingerNet focuses solely on the hand approaching and releasing motions, we only use the corresponding portions of the dataset. Specifically, we extract one second of motion before the grasp begins and after it ends as training data, discarding the rest.
We adopt a mirroring strategy similar to [70] to handle both hands simultaneously. During training, left-hand data are mirrored and combined with the right-hand data so that the model only needs to predict motions for the right hand. At inference time, we mirror the left-hand input, predict the corresponding motions using the model, and then mirror the output back to obtain the left-hand motions.
B.5 Smooth Transition between Navigation and Interaction
The navigation and interaction modules alternate to generate a long-horizon interaction sequence. Specifically, when the human is far from the object, the navigation module is activated to guide the human towards it. Once the human is within a threshold distance (set at 1 meter), the system switches to the interaction module to begin object manipulation. After the interaction is completed and the object is put down, control reverts to the navigation module to move towards the next object.
Since both modules use the initial human pose as an input condition, we ensure a smooth transition by feeding the last pose from the previous module into the next. We apply the same smoothing technique described in Sec. B.3 at the switching boundary to further enhance continuity.
C Implementation Details of Physics Tracker
We use IsaacGym [33] as the physics engine to perform simulation. The simulated character is skeleton-driven where each joint is modeled as a motor controlled through a PD controller. The control policy works at 30Hz, which is the same with generated kinematics motions, while the simulation runs at 120Hz. Our character has $62$ body links and $49$ controllable joints. This results in a state space $\mathbf{s}_{t}\in\mathbb{R}^{(62+|\mathcal{O}|)\times 13}$ including the position, orientation and linear and angular velocities of the character’s body links and the manipulating objects $\mathcal{O}$ depending on the given tracking motions. To track the target motion, additional observation $\mathbf{o}_{t}\in\mathbb{R}^{(62+|\mathcal{O}|)\times 7}$ including the target position and orientation of each body link and object is also introduced as the input to the control policy. The action space is $\mathbf{a}_{t}\in\mathbb{R}^{49\times 3}$ given each controllable joint has 3 degrees of freedom. All our control policies are optimized using PPO [45] as the backbone reinforcement learning algorithm with network parameters updated by Adam optimizers [24]. The hyperparameters used for policy training are listed in Tab. S2.
C.1 Reward Design
Since fingers take $30$ out of $49$ controllable joints and have a higher requirement in control precision for accurate object manipulation, we define a reward function containing three terms to evaluate the full-body pose and finger pose separately:
$r=0.8r_{\text{body}}+0.2r_{\text{hand}}+0.05r_{\text{energy}},$ (S1)
where $r_{\text{body}}$ evaluates the full-body poses excluding fingers, $r_{\text{hand}}$ evaluates the hand poses related to the target manipulating object, and $r_{\text{energy}}$ is an additional energy-related term penalizing end effectors’ abrupt movement. Note that in our implementation, at any time step, at most one object is activated as the manipulation target, while multiple objects may exist in the scene and would be manipulated one by one.
Body Part    Value
root (pelvis) $w_{q}=1,w_{p}=1$
lower abdomen $w_{q}=0.2$
upper abdomen $w_{q}=0.2$
chest $w_{q}=0.2$
neck $w_{q}=0.2$
head $w_{q}=0.2$
clavicles $w_{q}=0.1$
upper arms $w_{q}=0.2$
lower arms $w_{q}=0.2$
wrists $w_{q}=0.3,w_{p}=0.3$
thighs $w_{q}=0.5$
calfs $w_{q}=0.3$
feet $w_{q}=0.2,w_{p}=0.1$
target object $w_{q}=1,w_{p}=1$
Table S1: Unnormalized weights of each body link of the character and the target object when computing the reward regarding body pose errors (cf. $r_{\text{body}}$ in Equation S2). All weights unlisted are zero. The weights are normalized such that $\sum_{b}w_{p,b}=1$ and $\sum_{b}w_{q,b}=1$ before computing $r_{\text{body}}$.
Parameter Value
policy network learning rate $5\times 10^{-6}$
critic network learning rate $1\times 10^{-4}$
reward discount factor ($\gamma$) $0.95$
GAE discount factor ($\lambda$) $0.95$
surrogate clip range ($\epsilon$) $0.2$
number of PPO workers
   (simulation environment instances) $2048$
PPO replay buffer size $2048\times 8$
PPO batch size $256$
PPO optimization epochs 55
Table S2: Hyperparameters of policy training for physics tracker.
For full-body pose evaluation, we consider the orientation and position accuracy of the character’s body links and the target object:
[…]
where $\mathcal{B}$ is the set of all body links of the character and the activated target object at the current time step, $\mathbf{e}_{q,b}$ is the orientation error between each body link $b$ of the simulated character and that in the tracking motion, $\mathbf{e}_{p,b}$ is the position error, and $w_{q,b}$ and $w_{p,b}$ are the associated weights. We measure $\mathbf{e}_{q,b}$ using the radian of the angle needed to rotate the link $b$ to the target orientation, while $\mathbf{e}_{p,b}$ is measured by the Euclidean distance from the current position of link $b$ to the target position. All weight values are listed in Tab. S1. For position tracking, we focus only on the position of the root link and end effectors to ensure that the trajectory of end effectors (foot and hand excluding fingers) matches the tracking target, while relying on orientation tracking to ensure the naturalness of full-body poses.
For hand pose evaluation, to ensure that the object would be manipulated by hands correctly, we consider the fingers’ relative position to the target manipulating object when the hand is close to the object, or their relative position to the wrist when the hand is far away from the target object:
[…]
where $\mathcal{F}$ is the set of fingers, $\mathbf{e}_{f,o}$ is the position error between each finger $f$ of the simulated character and the tracking target position in the object’s local system, $\mathbf{e}_{f,w}$ is the position error in the wrist’s local system, and $\alpha$ is an interpolation coefficient depending on the distance between the hand and object in the generated tracking motions. We define $\alpha=1$ when the hand is equal to or less than 0.25m from the target object and $\alpha=0$ if the hand is at least 1m far away from the object in the tracking motion.
$r_{\text{energy}}$ takes into account the character’s end effectors to prevent jittering or abrupt movement with larger acceleration:
[…]
where $\mathcal{E}$ is the set of the character’s key end effectors (feet and hands excluding fingers), and $\mathbf{a}_{e}$ is the linear acceleration of the end effector $e$.
In contrast to recent works [3, 55, 32] on object manipulation, our reward function does not rely on the contact states between the hands (or fingers) and objects. This is because the contact data derived from the kinematic motions generated by diffusion models can be unreliable due to artifacts (cf. Fig. 7). Moreover, this approach aligns with the common scenario in which detailed contact information is typically absent in motions obtained from motion capture and video demonstrations [41].
C.2 Importance Sampling Strategy
Given that the generated kinematic motions can span several minutes and involve interactions with multiple objects, it would be highly inefficient to perform motion tracking from the very beginning of the sequence. Since locomotion without object manipulation and the phase following object grasping are relatively easier to learn, we adopt an importance sampling strategy that focuses on pre-grasp states to facilitate policy training. Specifically, we partition the parallel simulated physics environments into $|\mathcal{O}|$ batches, with each batch assigned to a specific manipulating object. In each batch, simulated characters are initialized in a random pre-grasp pose while interacting with the assigned object. This approach increases the likelihood that the policy learns the transition from the pre-grasp to the grasping phase, and subsequently, the locomotion that follows a successful grasp. Moreover, this strategy ensures that the policy consistently learns to interact with multiple objects simultaneously, thereby avoiding potential local minima during policy exploration.
D Inference Speed
Achieving real-time performance remains an open challenge for both diffusion-based and optimization-based methods. In our low-level motion generator, producing a 4-second motion sequence takes about four seconds for each diffusion model in stages 1, 3, and 4, while the optimization process in stage 2 requires roughly three minutes. Additionally, training the physics tracker on a 4-second interaction segment takes approximately 2 hours. All experiments were run on a single NVIDIA RTX 4000 GPU.
Figure S5: PhysHOI [55] fails to track kinematic motion, while our policy successfully drives the character to interact with diverse objects.
E Additional Experiemental Analysis
E.1 Implementation Details of Baseline
Since no prior work addresses the task of generating full-body motion, finger motion, and object motion from language instructions, we adapt prior works CHOIS [25] and GRIP [48] as fair baselines to compare against our system. CHOIS generates full-body motion and object motion from text and initial states, which also serves as our Stage 1. The output of CHOIS will serve as input to GRIP. GRIP consists of an arm denoising model and a two-stage finger motion generation model. It takes full-body motion (without finger motion) and object motion as input to denoise arm motion. It then proposes a hand sensor to extract hand-object spatial features and predict initial finger poses. In the second stage of finger motion generation, GRIP leverages these predicted finger poses to compute proximity features, which serve as input to refine the finger motion.
E.2 Results of Navigation Module
We report the evaluation scores for the navigation module in Tab. S3, using the metrics introduced in Sec. 7.2. The module is evaluated on the test split of the HumanML3D dataset [12]. It demonstrates that the module closely adheres to the input waypoints, with a relatively small joint position error when compared to the ground truth data, while maintaining reasonable foot sliding and foot height.
$T_{xy}$ $H_{\text{feet}}$ FSFS MPJPE Ours 4.00 1.48 0.58 10.87
Table S3: Navigation synthesis on the HumanML3D dataset [12]. The joint error is relatively small, demonstrating that the model aligns well with the input waypoints.
E.3 Results of Physics Tracker
We compare our physics tracker against PhysHOI [55], which trains reinforcement learning policies to control a humanoid character playing basketball. For our task, we have PhysHOI track the kinematic motion generated by our low-level motion generator. As shown in Fig. S5, PhysHOI fails to grasp objects, whereas our method successfully completes the task.
E.4 Human Perception Study Details
We show the interface of the human perceptual study for the high-level planner in Figure S6 and the low-level motion generator in Figure S7.
Figure S6: Interface of human perceptual study for the high-level planner.
Figure S7: Interface of human perceptual study for the low-level motion generator.